OpenStudy (anonymous):

degree 3 polynomial. It goes through the point (5,−2.1). Find the polynomial.

OpenStudy (freckles):

You are given f(x)=ax^3+bx^2+cx+d and f(5)=-2.1 . I don't think that is enough info to find all the constants there... Maybe it just wants to find a polynomial satisfying the above info ... so maybe you could just look at f(x)=ax^3+1 or f(x)=x^3+d etc...